How much of a 75% copper alloy should be melted into 62 kg of a 35% copper alloy to produce an alloy which is 50% copper?

Answer by checkley79(3341)

.75x+.35*62=.50(62+x)
.75x+21.70=31+.50x
.75x-50x=31-21.70
.25x=9.30
x=9.30/.25
x=37.2 kg of the 75% alloy is needed.
Proof:
.75*37.2+.35*62=.50(62+37.2)
27.9+21.70=.50*99.2
49.60=49.60